# Reasons to Reject? Aligning Language Models with Judgments

### v2

@@ Abstract
As humans, we consistently interact with our peers and receive feedback in the form of natural language. This language feedback allows us to maintain appropriate behavior, and rectify potential errors. The question arises naturally: can we use language feedback to align large language models (LLMs)? In contrast to previous research that aligns LLMs with scalar rewards, we present the first systematic exploration of alignment through the lens of language feedback (i.e., judgment). We start with an in-depth investigation of potential methods that can be adapted for aligning LLMs with judgments, revealing that these methods cannot fully capitalize on judgments. To facilitate more effective utilization of judgments, we propose a novel framework, Contrastive Unlikelihood Training (CUT), that allows for fine-grained inappropriate content detection and correction based on judgments. Our results show that, with merely 1317 off-the-shelf judgment data, CUT (LLaMA2-13b) can beat the 175B DaVinci003 and surpass the best baseline by 48.51 points on AlpacaEval. CUT (LLaMA2-chat-13b) can also align LLMs in an iterative fashion using up-to-date model-specific judgments, improving performance from 81.09 to 91.68 points on AlpacaEval. Further analysis suggests that judgments hold greater potential than rewards in LLM alignment.
@@ 1 Introduction
Large language models (LLMs) acquire extensive knowledge and remarkable reasoning capabilities through large-scale pre-training (Brown et al., 2020; Du et al., 2022; Touvron et al., 2023). To unleash the power of pre-trained LLMs for addressing real-world applications, it is crucial to ensure LLMs can follow human values (Ouyang et al., 2022). This process, known as alignment, has the potential to pave the way for a future in which artificial intelligence (AI) serves as a helpful and reliable ally for humanity (Wang et al., 2023b).
Figure 1 shows three typical paradigms to achieve alignment. The most straightforward one is learning from demonstrations, wherein demonstrations of desired responses to a set of instructions are collected to fine-tune LLMs (Wei et al., 2022; Ouyang et al., 2022). However, the performance gains diminish rapidly when scaling up the data size (Zhou et al., 2023). On the other hand, learning from feedback offers a more scalable approach (Ouyang et al., 2022; Bai et al., 2022a). One key advantage of feedback over demonstrations is that feedback can convey both positive and negative aspects, enabling the model to discern desirable and undesirable outcomes. In addition, feedback is tailored to the current model, adhering to the principle of teaching according to the learner’s aptitude.
Prior research on learning from feedback primarily focuses on value feedback (i.e., scalar rewards). Reinforcement learning (RL) techniques, particularly PPO algorithm (Schulman et al., 2017), are employed to optimize an LLM to maximize the rewards of its generated responses. However, PPO is known to be complex and often unstable (Zheng et al., 2023), which has prompted numerous efforts to simplify or stabilize the training process (Ramamurthy et al., 2023; Peng et al., 2023b; Dong et al., 2023). Another strand of work, named Hindsight (Zhang et al., 2023; Liu et al., 2023a), transforms scalar rewards to language instructions and employs supervised training on the updated data.
@@
Language feedback (i.e., judgment) is another kind of feedback that offers nuanced commendations and critiques through natural language expressions. Unlike scalar rewards, which are information-sparse for solely indicating the goodness of a response, judgments can elucidate the specific aspects that are good or bad, the rationale behind their evaluation, and suggestions for improvement. The above advantages suggest that aligning LLMs with judgments can be more beneficial (Saunders et al., 2022). Current approaches merely use judgments to prompt LLMs for an improved response (Bai et al., 2022b; Scheurer et al., 2022; Yu et al., 2023). But this leaves the judgment minimally involved in the subsequent alignment stage. This indirect utilization of judgments suffers from the incapability to learn from mistakes, which is the core spirit of learning from feedback.
In this study, we present an extensive investigation of potential methods that can be adapted for aligning LLMs with judgments. To facilitate a comprehensive aligning process, we propose a novel framework, Contrastive Unlikelihood Training (CUT), that enables fine-grained inappropriate content detection and correction based on judgments. CUT detects inappropriate content in a response by contrasting its generation probabilities under aligned and misaligned conditions and further penalizes the inappropriate content with unlikelihood training (UT) Welleck et al. (2020).
We carry out experiments for both offline and online alignment, wherein the target LLM learns from the off-the-shelf judgments and the judgments derived from self-generated responses, respectively. Extensive results on offline alignment demonstrate the effectiveness of CUT in learning from judgments in both cold-start (using unaligned base LLMs such as LLaMA2) and warm-start (using aligned base LLMs such as LLaMA2-chat) scenarios. Notably, when trained with only 1317 offline judgment data, CUT (LLaMA2-13b) attains a winning rate of 61.06 and outperforms the best baseline by 48.51 points on AlpacaEval. Furthermore, our online alignment experiments show that CUT is capable of iteratively refining LLMs (LLaMA2-chat-13b) using model-specific judgments, with a steady performance improvement from 81.09 to 91.68 points on AlpacaEval. Our analysis comparing rewards and judgments suggests that aligning LLMs with judgments offers significant potential and warrants future research.
@@ 2 Related Work
Prompting. Prompting does not alter the parameters of LLMs. Instead, it leverages judgments on previous responses to elicit improved responses from LLMs (Welleck et al., 2022; Akyurek et al., 2023). Judgments can be sourced from diverse aspects (Nathani et al., 2023) and the refinement process can be iterated multiple times (Yang et al., 2022; Peng et al., 2023a; Madaan et al., 2023). However, these methods consume more computation than single-pass generation and usually rely on the in-context learning capabilities of the LLMs (Brown et al., 2020; Liu et al., 2023b).
Fine-tuning. Fine-tuning aims to train a better LLM directly. Scalar rewards have been extensively used through RL, particularly PPO (Schulman et al., 2017; Ziegler et al., 2019; Yang et al., 2023). However, these RL approaches are notoriously unstable and complex (Zheng et al., 2023). To stabilize RL, Ramamurthy et al. (2023) propose to reduce the action space through truncation and Peng et al. (2023b) employ an advantage model and selective rehearsal. In addition, many studies aim to design simpler alternatives to RL. Dong et al. (2023); Touvron et al. (2023) treat rewards as a ranking criterion and simply train models using the best model-generated responses. There are also attempts to leverage the results of prompting for training a better model. That is, the improved response elicited by the judgment is employed as new training data (Scheurer et al., 2022, 2023; Yu et al., 2023). However, these methods still suffer from the incapability to learn from mistakes. Rafailov et al. (2023); Yuan et al. (2023); Song et al. (2023); Hong et al. (2024) demonstrate that LLMs themselves can be used as reward models and derive different training objectives to eliminate the need for RL. Zhang et al. (2023); Liu et al. (2023a) relabel the input using the reward received by the response, referred to as Hindsight. Hindsight allows LLMs to generate responses of different qualities.
@@ 3.2 Potential Solutions
The improved response $\hat{\bm{y}}$ can be directly paired with the initial instruction $\bm{x}$ to fine-tune the LLM under the MLE objective (Bai et al., 2022b; Scheurer et al., 2022, 2023).
$L_{i}=-\frac{1}{N}\sum_{t}\log p(\hat{y}_{t}|\hat{y}_{<t},\bm{x})$ (3)
@@ 4.3 Relation to Prior Solutions
Hindsight learns to generate responses of different qualities at the risk of increasing the likelihood of unsatisfactory responses. In comparison to Hindsight, CUT mitigates this issue by incorporating UT objectives for inappropriate tokens.
@@ 5 Experiments
LLaMA2-chat Base - 58.02 79.89 54.52 45.44 59.47 81.09
@@
LLaMA2-chat Base 29.21 15.00 22.78 23.44
@@
We implement CUT in two alignment settings, namely, online alignment and offline alignment, to demonstrate the overall effectiveness of CUT. Subsequently, to highlight the immense potential of aligning LLMs with judgments, we establish a comparison between learning from rewards and learning from judgments. The details of the model implementations are provided in Appendix A.1.
@@ 5.1 Offline Alignment
The offline setting utilizes off-the-shelf instruction-response-judgment triplets for alignment. This aims to check the feasibility of the CUT in learning from judgments prior to initiating the costly process of model-specific judgment annotation.
Tasks. We experiment on a general instruction-following task, and a specific NLP task (summarization). For Instruction following, we train models with 1317 examples from Shepherd (Wang et al., 2023a) and evaluate models on four ranking-based and one generation-based LLM benchmarks. The ranking benchmarks are 25-shot ARC, 10-shot HellaSwag, 5-shot MMLU, and 0-shot TruthfulQA. The generation benchmark is AlpacaEval, where GPT4 judges the winning rate of the responses generated by our models against DaVinci003. For Summarization, we use the dataset with judgment annotations from Saunders et al. (2022). We train models on 10827 training examples and report ROUGE scores (Lin, 2004) on 1939 test examples. See Appendix A.3 for more details.
Baselines. We experiment with two base models, LLaMA2-13b and LLaMA2-chat-13b, to demonstrate the efficacy of CUT in both cold-start and warm-start scenarios, respectively. The baselines include the base model without further fine-tuning, and the three groups of judgment-based alignment methods: 1 The base form of Forward Prediction described in Eq. 1 (Weston, 2016) and Mistake Analysis that uses self-generated judgments and unguided fine-tuning strategy (Chen et al., 2024); (2) The base form of Hindsight described in Eq. 4 (Zhang et al., 2023); (3) The base form of ILF described in Eq. 3 (Scheurer et al., 2022) and CDF (Yu et al., 2023) that change the learning objective from MLE to Direct Preference Optimization (DPO) (Rafailov et al., 2023).
@@
Results. The results of the general instruction-following and summarization are presented in Table 2 and 3, respectively. For cold-start scenarios (LLaMA2 as the base model), CUT improves the winning rate on AlpacaEval from 1.87 to 61.06, where CUT beats the 175B DaVinci003 and surpasses the best baseline (Mistake Analysis) by 48.51 points. Moreover, CUT improves the base model by 13.08 points on TruthfulQA. This implies that CUT can effectively mitigate hallucinations. Conversely, most baselines improve marginally or experience performance drops on TruthfulQA. This is likely due to their application of the MLE objective on error-prone responses, which reduces factuality in response generation. In terms of ARC, HellaSwag, and MMLU, CUT remains competitive with the base model, indicating CUT suffers less from the alignment tax problem (Ouyang et al., 2022). For single NLP task (i.e., summarization) experiments, CUT surpasses the best baseline (i.e., Forward Prediction) by 1.38 rougeLsum scores. Overall, the results show that CUT is effective in transforming LLMs into both performant generalist and specialist models.
@@
The performance improvements of warm-start scenarios (LLaMA2-chat as the base model) are consistent with the cold-start ones, showcasing the efficacy of CUT in learning from judgments in both cold-start and warm-start scenarios. The two ILF methods (ILF and CDF) outperform methods from Forward Prediction and Hindsight groups on AlpacaEval in warm-start but perform worse in cold-start scenarios. This may be due to that ILF methods heavily relies on the base model in producing high-quality improved responses, making it less effective in cold-start scenarios.
@@
Model Generalist Specialist
LLaMA2-chat 45.44 23.44
CUT 52.58 40.12
- $L_{1}$ 39.01 37.46
- first part of $L_{2}$ 26.37
- second part of $L_{2}$ 47.24 33.92
- Inappropriate Token Detection 0 0
- Dynamic Weighting 48.84 40.05
Framework Design. To investigate the effectiveness of two contrasts employed by CUT, we perform ablation studies by eliminating certain training signals. The results are shown in Table 5. Removing the contrast between Align-N and Misalign (- $L_{1}$) substantially reduces the performance of TruthfulQA. This finding highlights that the UT objective plays a crucial role in mitigating hallucinations. The exclusion of the contrast between Align-P and Align-N can be implemented in two ways. We can either remove the first part or the second part of $L_{2}$. As seen, the impact of removing Align-P is more pronounced than removing Align-N on the summarization task. This may be attributed to the necessity of positive examples for adapting the LLM to a specific task. Furthermore, we introduce an additional ablated variant in which the inappropriate token detection (Eq. 5) is omitted (- Inappropriate Token Detection). Concretely, we simply apply UT for all tokens in misaligned responses instead. Intriguingly, we find that this approach fails to converge during training. This observation underscores the importance of inappropriate token detection. Lastly, removing the dynamic weighting term ($p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})^{\gamma}$ in Eq. 6) also impacts the effectiveness of the CUT framework, particularly in general instruction-following tasks.
@@ 5.2 Online Alignment
In this section, we move to a more pragmatic scenario where the target LLM directly learns from the judgments associated with its own responses.
5.2.1 Iterative Alignment
Setup. As mentioned in § 3.1, the online alignment process can be conducted iteratively, akin to how humans continuously refine their behaviors through ongoing feedback. Specifically, we apply the following three steps repeatedly:
@@ 5.2.1 Iterative Alignment
where the target LLM is LLaMA2-chat. In each iteration, we sample 1000 distinct instructions from Stanford Alpaca (Taori et al., 2023). We ask GPT4 for drafting judgments, which has been proven to produce high-quality annotations (Cui et al., 2023). Annotation details are elaborated in Appendix A.2. Note that most responses from LLaMA2-chat receive positive judgments, resulting in a large proportion of Align-P examples. We found downsampling Align-P examples is beneficial to the online alignment (see Appendix A.4). We evaluate models on ARC, HellaSwag, MMLU, TruthfulQA, and AlpacaEval.

### v3

@@ Abstract
As humans, we consistently interact with our peers and receive feedback in the form of natural language. This language feedback allows us to maintain appropriate behavior, and rectify potential errors. The question arises naturally: can we use language feedback to align large language models (LLMs)? In contrast to previous research that aligns LLMs with scalar rewards, we present the first systematic exploration of alignment through the lens of language feedback (i.e., judgment). We start with an in-depth investigation of potential methods that can be adapted for aligning LLMs with judgments, revealing that these methods cannot fully capitalize on judgments. To facilitate more effective utilization of judgments, we propose a novel framework, Contrastive Unlikelihood Training (CUT), that allows for fine-grained inappropriate content detection and correction based on judgments. Our results show that, with merely 1317 off-the-shelf judgment data, CUT can beat the 175B DaVinci003 and surpass the best baseline by 50.84 points on AlpacaEval using LLaMA2-13b. CUT can also align LLMs in an iterative fashion using up-to-date model-specific judgments, improving performance from 81.09 to 91.68 points on AlpacaEval using LLaMA2-chat-13b. Further analysis suggests that judgments hold greater potential in LLM alignment than rewards.
@@ 1 Introduction
Large language models (LLMs) acquire substantial world knowledge and reasoning capabilities through large-scale pre-training (Brown et al., 2020; Du et al., 2022; Touvron et al., 2023). To unleash the power of pre-trained LLMs for real-world applications, it is crucial to ensure that LLMs can follow human preferences and values (Ouyang et al., 2022). This process, known as alignment, is critical for making artificial intelligence a helpful and reliable ally for humanity (Wang et al., 2023b).
Figure 1 illustrates three paradigms to achieve alignment. The most straightforward one is learning from demonstrations, wherein demonstrations of desired responses to a set of instructions are collected and used to fine-tune LLMs in a supervised fashion (Wei et al., 2022; Ouyang et al., 2022). However, the performance gains diminish rapidly as the data size increases (Zhou et al., 2023; Fu et al., 2024). In contrast, learning from feedback (rewards or judgements) offers a more scalable approach (Ouyang et al., 2022; Bai et al., 2022a). One significant advantage of feedback over demonstrations is that feedback can convey both positive and negative aspects, enabling the model to discern desirable and undesirable outcomes. In addition, feedback is tailored to the current model, adhering to the principle of teaching according to the learner’s aptitude.
Prior research on learning from feedback primarily focuses on value feedback (i.e., scalar rewards), employing reinforcement learning (RL) algorithms, such as PPO (Schulman et al., 2017), to optimize an LLM to maximize the rewards of its generated responses. Nevertheless, scalar rewards are information-sparse for solely indicating the goodness of a response. On the other hand, language feedback (i.e., judgment) can offer more nuanced commendations and critiques through natural language expressions (Saunders et al., 2022). Specifically, judgments can elucidate the specific aspects that are good or bad, the rationale behind their evaluation, and suggestions for improvement. The above suggests that aligning LLMs with judgments can be more advantageous.
@@
In this study, we present an extensive investigation of potential methods that can be adapted for aligning LLMs with judgments. To facilitate a comprehensive aligning process, we propose a novel framework, Contrastive Unlikelihood Training (CUT), that enables fine-grained inappropriate content detection and correction based on judgments. CUT detects inappropriate content in a response by contrasting its generation probabilities under aligned and misaligned conditions and further penalizes the inappropriate content with unlikelihood training Welleck et al. (2020).
We carry out experiments for both offline and online alignment, wherein the target LLM learns from the off-the-shelf judgments and the judgments derived from self-generated responses, respectively. Extensive results on offline alignment demonstrate the effectiveness of CUT in learning from judgments in both cold-start (using unaligned base LLMs such as LLaMA2) and warm-start (using aligned base LLMs such as LLaMA2-chat) scenarios. Notably, when trained with only 1317 offline judgment data, CUT attains a winning rate of 61.06 and outperforms the best baseline by 50.84 points on AlpacaEval using LLaMA2-13b. Furthermore, our online alignment experiments show that CUT is capable of iteratively refining LLMs using model-specific judgments, with a steady performance improvement from 81.09 to 91.68 points on AlpacaEval using LLaMA2-chat-13b. Our analysis comparing rewards and judgments suggests that aligning LLMs with judgments offers significant potential and warrants future research.
@@ 2 Related Work
Prompting. Prompting does not alter the parameters of LLMs. Instead, it leverages judgments on previous responses to elicit improved responses from LLMs (Welleck et al., 2022; Akyurek et al., 2023). Judgments can be sourced from diverse aspects (Nathani et al., 2023) and the refinement process can be iterated multiple times (Yang et al., 2022; Peng et al., 2023a; Madaan et al., 2023). However, these methods rely on the in-context learning capabilities of the LLMs and consume more computation than single-pass generation (Brown et al., 2020; Liu et al., 2023b).
Fine-tuning. Fine-tuning aims to train an LLM that can generate better responses immediately. Scalar rewards have been extensively used through the lens of RL, particularly PPO (Schulman et al., 2017; Ziegler et al., 2019; Yang et al., 2023). However, PPO is known to be complex and unstable (Zheng et al., 2023), which has attracted numerous efforts to simplify or stabilize the training process (Ramamurthy et al., 2023; Peng et al., 2023b; Dong et al., 2023; Touvron et al., 2023; Rafailov et al., 2023; Yuan et al., 2023; Song et al., 2023; Hong et al., 2024). Another strand of work, named Hindsight (Zhang et al., 2023; Liu et al., 2023a), transforms scalar rewards to language instructions and teach LLMs to generate responses of different qualities. There are also attempts to leverage the results of prompting for training a better model. That is, the improved response elicited by the judgment is employed as new training data (Scheurer et al., 2022, 2023; Yu et al., 2023). However, these methods still suffer from the incapability to learn from mistakes, which is the core spirit of learning from feedback.
@@ 3.2 Potential Solutions
  • ILF-MLE: The improved response $\hat{\bm{y}}$ can be directly paired with the initial instruction $\bm{x}$ to fine-tune the LLM under the MLE objective (Bai et al., 2022b; Scheurer et al., 2022, 2023).
    $L_{i}^{mle}=-\frac{1}{N}\sum_{t}\log p(\hat{y}_{t}|\hat{y}_{<t},\bm{x})$ (3)
  • ILF-DPO: Yu et al. (2023) demonstrates that the improved response $\hat{\bm{y}}$ and the original response $\bm{y}$ can be used jointly as a pairwise comparison, where $\hat{\bm{y}}$ is a more preferred response to $x$ compared to $\bm{y}$. As a result, preference learning algorithms, such as direct preference optimization (DPO) (Rafailov et al., 2023), can be adopted to fine-tune the LLM: $L_{i}^{dpo}=\textbf{DPO}(\bm{x},\bm{y},\hat{\bm{y}})$.
@@ 4.3 Relation to Prior Solutions
Hindsight learns to generate responses of different qualities at the risk of increasing the likelihood of unsatisfactory responses. In comparison to Hindsight, CUT mitigates this issue by incorporating both likelihood and unlikelihood training objectives.
@@ 5 Experiments
LLaMA2-chat-13b Base - 58.02 79.89 54.52 45.44 59.47 81.09
@@
LLaMA2-chat-13b Base 29.21 15.00 22.78 23.44
@@
To provide a comprehensive assessment of CUT, we implement it in two alignment scenarios: offline alignment and online alignment. In the offline alignment experiments, we perform extensive analysis on the adaptability and universality of CUT across different model and task configurations. In the online alignment experiments, we additionally explore the possibility of building an automatic judgment model. Lastly, to highlight the potential of aligning LLMs with judgments, we establish a comparison between learning from rewards and learning from judgments.
Tasks. We experiment on both general instruction-following and a specific NLP task (summarization). For Instruction following, we evaluate models on both AlpacaEval and four additional conventional NLP benchmarks: 25-shot ARC, 10-shot HellaSwag, 5-shot MMLU, and 0-shot TruthfulQA. For AlpacaEval, we report the winning rate of the responses generated by our models against DaVinci003 using GPT4 as the judge. The four conventional NLP benchmarks are ranking-based and we report accuracies. For Summarization, we use the dataset from Saunders et al. (2022) and report ROUGE scores (Lin, 2004) on 1939 test examples. See Appendix A.3 for more details.
Baselines. The baselines include the base model without further fine-tuning, and the three groups of judgment-based alignment methods: (1) The Forward Prediction method described in Eq. 1 (Weston, 2016); (2) The Hindsight method described in Eq. 4 (Zhang et al., 2023); (3) ILF-MLE described in Eq. 3 (Scheurer et al., 2022), and ILF-DPO (Yu et al., 2023) that change the learning objective from MLE to DPO. The details of the model implementations are provided in Appendix A.1.
@@ 5.1 Offline Alignment
The offline setting utilizes off-the-shelf instruction-response-judgment triplets for alignment. This aims to validate and analyze CUT in controlled environments prior to initiating the costly process of model-specific judgment annotation. For instruction following, we train models with 1317 examples from Shepherd (Wang et al., 2023a). For summarization, we use the 10827 training examples with judgment annotations from Saunders et al. (2022).
@@
Results. The results of the general instruction-following and summarization are presented in Table 2 and 3, respectively. For cold-start scenarios (LLaMA2-13b as the base model), CUT improves the winning rate on AlpacaEval from 1.87 to 61.06, where CUT beats the 175B DaVinci003 and surpasses the best baseline (Hindsight) by 50.84 points. Moreover, CUT improves the base model by 13.05 points on TruthfulQA. This implies that CUT can effectively mitigate hallucinations. Conversely, most baselines improve marginally or experience performance drops on TruthfulQA. This is likely due to their application of the MLE objective on error-prone responses, which reduces factuality in response generation. In terms of ARC, HellaSwag, and MMLU, CUT remains competitive with the base model, indicating CUT suffers less from the alignment tax problem (Ouyang et al., 2022). For single NLP task (i.e., summarization) experiments, CUT surpasses the best baseline (i.e., Forward Prediction) by 1.38 rougeLsum scores. Overall, the results show that CUT is effective in transforming LLMs into both performant generalist and specialist models.
@@
The performance superiority of CUT in warm-start scenarios (LLaMA2-chat-13b as the base model) are consistent with the cold-start ones. The two ILF methods (ILF-MLE and ILF-DPO) outperform methods from Forward Prediction and Hindsight groups on AlpacaEval in warm-start but perform worse in cold-start scenarios. This may be due to that ILF methods heavily rely on the base model in producing high-quality improved responses, making it less effective in cold-start scenarios.
Model Generalist Specialist
LLaMA2-chat-13b 45.44 23.44
CUT 52.58 40.12
- $L_{1}$ 39.01 37.46
- first part of $L_{2}$ 26.37
- second part of $L_{2}$ 47.24 33.92
- Inappropriate Token Detection 0 0
- Dynamic Weighting 48.84 40.05
Ablation Study. To investigate the effectiveness of two contrasts employed by CUT, we perform ablation studies by eliminating certain training signals. The results are shown in Table 4. Removing the contrast between Align-N and Misalign (- $L_{1}$) substantially reduces the performance of TruthfulQA. This finding highlights that the UT objective plays a crucial role in mitigating hallucinations. The exclusion of the contrast between Align-P and Align-N can be implemented in two ways. We can either remove the first part or the second part of $L_{2}$. As seen, the impact of removing Align-P is more pronounced than removing Align-N on the summarization task. This may be attributed to the necessity of positive examples for adapting the LLM to a specific task. Furthermore, we introduce an additional ablated variant in which the inappropriate token detection (Eq. 5) is omitted (- Inappropriate Token Detection). Concretely, we simply apply UT for all tokens in misaligned responses instead. Intriguingly, we find that this approach fails to converge during training. This observation underscores the importance of inappropriate token detection. Lastly, removing the dynamic weighting term ($p(y_{t}|\bm{y}_{<t},\bm{x},\bm{j}^{-})^{\gamma}$ in Eq. 6) also impacts the effectiveness of CUT, particularly in general instruction-following tasks.
@@ 5.2 Online Alignment
In this section, we move to a more pragmatic scenario where the target LLM directly learns from the judgments associated with its own responses. As mentioned in § 3.1, the online alignment process can be conducted iteratively, akin to how humans continuously refine their behaviors through ongoing feedback. Specifically, we apply the following three steps repeatedly:
@@
where the target LLM is LLaMA2-chat-13b. In each iteration, we sample 1000 distinct instructions from Stanford Alpaca (Taori et al., 2023). We ask GPT4 for drafting judgments, which has been proven to produce high-quality annotations (Cui et al., 2023). Annotation details are elaborated in Appendix A.2. Note that most responses from LLaMA2-chat-13b receive positive judgments, resulting in a large proportion of Align-P examples. We found downsampling Align-P examples is beneficial to the online alignment (see Appendix A.4). We evaluate models on ARC, HellaSwag, MMLU, TruthfulQA, and AlpacaEval.